an electron released from rest in uniform electric field of magnitude e=220 n/c and gains speed. find magnitude of force applied to it


1
Expert's answer
2021-10-10T16:02:34-0400

The field of magnitude "E" acts on the point charge "q" with force "F = q E". Therefore, "F = 1.6 \cdot 10^{-19} C\cdot 220 \frac{N}{C} \approx 3.5 \cdot 10^{-17} N".